Dublin University Rifle Club Dublin University Rifle Club

c/o DUCAC Trinity College Dublin

Dublin 2 Ireland

Removal Order Application

I, hereby apply for a Removal Order to

authorise of D.U.R.C.

to use/have in his/her possession the following Rifles and Ammunition

at

on (dates)

.22 calibre Target Rifles

Walther KK200  011301
Anschutz 1913  1 179384x
2 175137x
4 1346313
7 135852
8 125738
9 129062
10 172895x
11 173313x

rounds of .22 calibre ammunition

.177 calibre Air Rifles

Feinwerkbau 601 Q  99983
R  24743
S 119439
T 109111
V  99752
Feinwerkbau 603  312234

air rifle pellets

To be collected by on Please Post.

Signed: